# BiomedCLIP: a multimodal biomedical foundation model pretrained from fifteen million scientific image-text pairs

### v2

@@ Abstract
Biomedical data is inherently multimodal, comprising physical measurements and natural language narratives. A generalist biomedical AI model needs to simultaneously process different modalities of data, including text and images. Therefore, training an effective generalist biomedical model requires high-quality multimodal data, such as parallel image-text pairs. Here, we present PMC-15M, a novel dataset that is two orders of magnitude larger than existing biomedical multimodal datasets such as MIMIC-CXR, and spans a diverse range of biomedical image types. PMC-15M contains 15 million biomedical image-text pairs collected from 4.4 million scientific articles. Based on PMC-15M, we have pretrained BiomedCLIP, a multimodal foundation model, with domain-specific adaptations tailored to biomedical vision-language processing. We conducted extensive experiments and ablation studies on standard biomedical imaging tasks from retrieval to classification to visual question-answering (VQA). BiomedCLIP achieved new state-of-the-art results in a wide range of standard datasets, substantially outperforming prior approaches. Intriguingly, by large-scale pretraining on diverse biomedical image types, BiomedCLIP even outperforms state-of-the-art radiology-specific models such as BioViL in radiology-specific tasks such as RSNA pneumonia detection. In summary, BiomedCLIP is a fully open-access foundation model that achieves state-of-the-art performance on various biomedical tasks, paving the way for transformative multimodal biomedical discovery and applications. We release our models at aka.ms/biomedclip to facilitate future research in multimodal biomedical AI.
@@ Introduction
Biomedical data is inherently multimodal, comprising both physical measurements and natural-language narratives [1, 2]. However, only a fraction of such data is interpreted and made available in structured forms, as manual curation is costly and hard to scale. Multimodal learning can alleviate the curation bottleneck by leveraging self-supervision from cross-modal correspondence. It also helps uncover predictive signals that may remain latent in standalone modalities through multimodal fusion. As a result, biomedical vision-language foundation models [3, 4, 5, 6] have been developed to collectively model biomedical images (e.g., digital pathology and radiology images) and biomedical text (e.g., scientific papers and clinical notes). There are two key advantages of biomedical vision-language foundation models. First, by pretraining on large-scale data using self-supervised learning, these models excel in low-resource settings, where only a small amount of fine-tuning data is accessible, across a diverse range of downstream applications, thereby substantially decreasing the cost associated with data annotation. Second, by pretraining on high-quality parallel image-text data, these models can solve tasks that require a comprehensive understanding of both image and text data, such as image description generation and visual question-answering.
@@
Based on PMC-15M, we have pretrained BiomedCLIP, a state-of-the-art biomedical vision-language foundation model that excels in a wide range of downstream applications such as cross-modal retrieval, zero-shot image classification, and medical visual question answering. BiomedCLIP outperformed the general-domain CLIP model [7] by a wide margin across the board, confirming the importance of exploring domain-specific models for biomedicine. By pretraining on much larger and more diverse data, BiomedCLIP also substantially outperformed prior state-of-the-art biomedical vision-language models such as PubMedCLIP and MedCLIP, especially in low-resource settings. Surprisingly, BiomedCLIP even outperformed radiology-specific state-of-the-art models such as BioViL [18] on radiology-specific tasks such as RSNA pneumonia detection [19], highlighting the potential of positive transfer in pretraining from diverse image categories. To facilitate future biomedical multimodal research, we will release our BiomedCLIP models and scripts to reproduce PMC-15M at aka.ms/biomedclip upon publication.
@@ Results
We next evaluate the performance of BiomedCLIP for biomedical image classification on five datasets (see Methods). We investigate the zero-shot performance on BiomedCLIP models, where a short text prompt describing the class is used to assist the classification (Supplementary Table 10). BiomedCLIP exhibits superior performance on zero-shot classification, attaining the highest overall accuracy (mean of the scores) across all the datasets (Fig. 3A,B). PubMedCLIP was continual pretrained using radiology data so it outperforms the general-domain CLIP on the RSNA benchmark but generally yields inferior performance on the digital pathology benchmarks such as LC25000 and TCGA-TIL. PLIP [6] was pretrained on digital pathology data from social media, so performs reasonably well on some of the pathology benchmarks, but yields much inferior performance on RSNA. Surprisingly, PLIP also performs rather poorly on the pathology benchmark PCam. We suspect that the pathology images in PCam (lymph nodes with metastatic tumors) might be under-represented in the social media data PLIP was pretrained on. In contrast, BiomedCLIP achieves good results on all datasets from diverse image sources, again indicating the importance of pretraining on a diverse and large-scale dataset for building a foundational vision-language model.
@@
Finally, we evaluate BiomedCLIP on medical visual question answering (VQA). Following the standard approach in prior work, we formulate VQA as a classification task using the METER [32] framework (see Methods). We evaluate the models on two standard datasets: VQ-RAD and SLAKE (Fig. 4A). We report the overall accuracy as well as respective accuracies for open-ended questions and closed-ended questions. Again, BiomedCLIP exhibits superior performance compared to the radiology-specific state-of-the-art PubMedCLIP on both radiology benchmarks, with about one point and six points increases in overall accuracy respectively. BiomedCLIP’s gains are particularly pronounced for open-ended questions in VQA-RAD and for all question types in SLAKE.
@@ Methods
We compare our method against four competing methods CLIP, MedCLIP, and PubMedCLIP. MedCLIP [26] extends the pretraining to include large unpaired images and texts through contrastive learning. It uses the pretrained BioClinicalBERT and Swin Transformer [48] as the backbone text encoder and visual encoder respectively, and fine-tunes on MIMIC-CXR and CheXpert datasets. PubMedCLIP [25] fine-tunes CLIP on the Radiology Objects in COntext (ROCO) dataset [13], which consists of 80K radiology image-text pairs drawn from PubMed articles. All the models are adapted into ELEVATER for evaluation.
@@
We utilize the METER [32] framework to facilitate our experiments on visual question answering (VQA). It formulates the VQA task as a classification task. The core module of METER is a transformer-based co-attention multimodal fusion module that produces cross-modal representations over the image and text encodings, which are then fed to a classifier for predicting the final answer. We compare BiomedCLIP with general-domain CLIP, MAML (Model-Agnostic Meta-Learning) network pretrained only on visual data, and the state-of-the-art PubMedCLIP. All three models were fine-tuned for VQA tasks using the QCR (Question answering via Conditional Reasoning) framework [49] that alternatively uses a MLP-based attention networks with conditional reasoning as the fusion module. Results for MAML, CLIP and PubMedCLIP are collected from [25].
We evaluated our method and competing methods on VQA-RAD [50]. VQA-RAD consists of $315$ radiology images and $3,515$ question-answer pairs that were manually constructed by clinicians. Images in the test set are also present in training set but the question-answer pairs do not overlap. SLAKE (English only) [51] consists of $642$ radiology images and over $7,000$ question-answer pairs annotated by experienced physicians. It covers more human body parts than VQA-RAD and does not have common images between the training and test sets. See Supplementary Table 1 for details.
@@
where $\tau$ is a learnable temperature parameter, directly optimized during training as a log-parameterized multiplicative scalar; ${\bm{I}}_{i}$ and ${\bm{T}}_{i}$ are embeddings for the $i$-th image and text, produced by a linear projection layer on top of the image encoder and text encoder. Rather than initializing with pretrained weights, CLIP trains the image encoder and text encoder from scratch. For the image encoder, CLIP considers two different architectures, ResNet-50 [52] and Vision Transformer [ViT; 27]. The text encoder is effectively GPT-2 [24] based on transformer [53].
@@
On the text side, we replace a blank-slate GPT-2 with a pretrained language model more suited for biomedicine. Specifically, we initialize with PubMedBERT, which shows substantial gains from domain-specific pretraining [20]. Correspondingly, for the tokenizer, we replace Byte-Pair Encoding [BPE; 54] with WordPiece [55], which uses unigram-based likelihood rather than shattering all words to characters and greedily forming larger tokens based on frequency. The original CLIP uses a context of 77 tokens, but biomedical text is typically longer, as shown in Fig. 1A. We thus increase the context size to 256, which covers 90% of PMC captions. Supplementary Table 1 shows that both modifications bring substantial improvements over the original CLIP model on the validation set.

### v3

@@ Abstract
Biomedical data is inherently multimodal, comprising physical measurements and natural language narratives. A generalist biomedical AI model needs to simultaneously process different modalities of data, including text and images. Therefore, training an effective generalist biomedical model requires high-quality multimodal data, such as parallel image-text pairs. Here, we present PMC-15M, a novel dataset that is two orders of magnitude larger than existing biomedical multimodal datasets such as MIMIC-CXR, and spans a diverse range of biomedical image types. PMC-15M contains 15 million biomedical image-text pairs collected from 4.4 million scientific articles. Based on PMC-15M, we have pretrained BiomedCLIP, a multimodal foundation model, with domain-specific adaptations tailored to biomedical vision-language processing. We conducted extensive experiments and ablation studies on standard biomedical imaging tasks from retrieval to classification to visual question-answering (VQA). BiomedCLIP achieved new state-of-the-art results in a wide range of standard datasets, including retrieval, image classification, and visual question answering, substantially outperforming prior approaches. Intriguingly, by large-scale pretraining on diverse biomedical image types, BiomedCLIP even outperforms state-of-the-art radiology-specific models such as BioViL in radiology-specific tasks such as RSNA pneumonia detection. We further illustrated how BiomedCLIP could promote privacy-preserving analysis of proprietary data by using PMC-15M as a proxy. In summary, BiomedCLIP is a fully open-access foundation model that achieves state-of-the-art performance on various biomedical tasks, paving the way for transformative multimodal biomedical discovery and applications. We release our models at aka.ms/biomedclip to facilitate future research in multimodal biomedical AI.
@@ Introduction
Biomedical data is inherently multimodal, comprising both physical measurements and natural-language narratives [1, 2]. However, only a fraction of such data is interpreted and made available in structured forms, as manual curation is costly and hard to scale. Multimodal learning can alleviate the curation bottleneck by leveraging self-supervision from cross-modal correspondence. It also helps uncover predictive signals that may remain latent in standalone modalities through multimodal fusion. As a result, biomedical vision-language foundation models [3, 4, 5, 6, 7, 8, 9, 10] have been developed to collectively model biomedical images (e.g., digital pathology and radiology images) and biomedical text (e.g., scientific papers and clinical notes). There are two key advantages of biomedical vision-language foundation models. First, by pretraining on large-scale data using self-supervised learning, these models excel in low-resource settings, where only a small amount of fine-tuning data is accessible, across a diverse range of downstream applications, thereby substantially decreasing the cost associated with data annotation. Second, by pretraining on high-quality parallel image-text data, these models can solve tasks that require a comprehensive understanding of both image and text data, such as image description generation and visual question-answering.
@@
Based on PMC-15M, we have pretrained BiomedCLIP, a state-of-the-art biomedical vision-language foundation model that excels in a wide range of downstream applications such as cross-modal retrieval, zero-shot image classification, and medical visual question answering. BiomedCLIP outperformed the general-domain CLIP model [11] by a wide margin across the board, confirming the importance of exploring domain-specific models for biomedicine. By pretraining on much larger and more diverse data, BiomedCLIP also substantially outperformed prior state-of-the-art biomedical vision-language models such as PubMedCLIP and MedCLIP, especially in low-resource settings. Surprisingly, BiomedCLIP even outperformed radiology-specific state-of-the-art models such as BioViL [22] on radiology-specific tasks such as RSNA pneumonia detection [23], highlighting the potential of positive transfer in pretraining from diverse image categories. Finally, we explore the possibility of using PMC-15M as a proxy for analyzing proprietary data using external models. To facilitate future biomedical multimodal research, we will release our BiomedCLIP models and scripts to reproduce PMC-15M at aka.ms/biomedclip upon publication.
@@ Results
We next evaluate the performance of BiomedCLIP for biomedical image classification on five datasets (see Methods). We investigate the zero-shot performance on BiomedCLIP models, where a short text prompt describing the class is used to assist the classification (Supplementary Table 10). BiomedCLIP exhibits superior performance on zero-shot classification, attaining the highest overall accuracy (mean of the scores) across all the datasets (Fig. 3A,B). PubMedCLIP was continual pretrained using radiology data so it outperforms the general-domain CLIP on the RSNA benchmark but generally yields inferior performance on the digital pathology benchmarks such as LC25000 and TCGA-TIL. PLIP [7] was pretrained on digital pathology data from social media, so performs reasonably well on some of the pathology benchmarks, but yields much inferior performance on RSNA. Surprisingly, PLIP also performs rather poorly on the pathology benchmark PCam. We suspect that the pathology images in PCam (lymph nodes with metastatic tumors) might be under-represented in the social media data PLIP was pretrained on. Med-Flamingo [10] was pre-trained as a few-shot learner on interleaved image-text data, with its pre-training data formulated as open-ended visual question answering (VQA). It almost always generates the same answer or choice regardless of the prompt template or test example, likely due to its pre-training nature, which limits its capability for zero-shot multimodal reasoning and hinders its ability to follow instructions for closed-set classification. In contrast, BiomedCLIP achieves good results on all datasets from diverse image sources, again indicating the importance of pretraining on a diverse and large-scale dataset for building a foundational vision-language model.
@@
After observing the superior performance of BiomedCLIP on image classification, we then evaluate it on medical visual question answering (VQA). Following the standard approach in prior work, we formulate VQA as a classification task using the METER [36] framework (see Methods). We evaluate the models on two standard datasets: VQ-RAD and SLAKE (Fig. 4A). We report the respective accuracies for open-ended questions and closed-ended questions as well as the token-level F1. Again, BiomedCLIP exhibits superior performance compared to the radiology-specific state-of-the-art PubMedCLIP on both radiology benchmarks, with about one point and six points increases in overall accuracy respectively. BiomedCLIP’s gains are particularly pronounced for open-ended questions in VQA-RAD and for all question types in SLAKE. Furthermore, we see that BiomedCLIP achieves comparable performance comparing to the much larger model Med-PalM M (562B) [37] and the instruction-aware fine-tuned model BiomedGPT-B [6]. Notably, LLaVA-Med [38], which integrates BiomedCLIP as the core vision encoder into the LLaVA framework and further fine-tunes it to follow open-ended instructions, achieves the highest VQA performance on both datasets.
@@
BiomedCLIP and PMC-15M enable privacy-preserving analysis of proprietary data
Finally, we investigate whether BiomedCLIP and PMC-15M can be used for privacy-preserving analysis of proprietary patient data (Fig. 5A). Proprietary patient data are often not allowed to be shared with external models, such as GPT4, preventing clinicians from using these models to analyze them. To tackle this, our idea is to use PMC-15M as a proxy for proprietary patient data. In particular, for each proprietary data point, we will use BiomedCLIP to retrieve the most similar data points in PMC-15M. We will then query external models using these similar public data points and aggregate the answers as the proxy output for the proprietary data point. During this process, the proprietary data point can be analyzed using external models without being exposed to GPT4.
To test the accuracy of this approximation, we collected $980$ radiology images and the associated clinical reports from Providence (see Methods) and treated them as proprietary patient data. We then compared the CheXbert-based labels assigned to the proprietary data and those labels assigned to the most similar data points in PMC-15M. We observe a good agreement between these labels in terms of recall and F1, which is substantially higher than the agreements by PLIP and CLIP. In particular, our method can get 88.80 and 95.04 recall on lung opacity and atelectasis (Fig. 5B), indicating that BiomedCLIP can accurately retrieve relevant data points in PMC-15M as high-quality proxies of proprietary radiology data. Meanwhile, BiomedCLIP gets a lower recall than CLIP on Cardiomegaly (Fig. 5B), suggesting that the approximation using BiomedCLIP could be limited by the captions in the scientific papers, which might not present all conditions as comprehensively as clinical reports. Finally, we illustrated an example proprietary image from Providence with the top three retrieved PMC images using BiomedCLIP (Fig. 5C), and observed that the captions of the retrieved PMC images could cover the main findings from the clinical report of the proprietary image, further raising our confidence that BiomedCLIP and BiomedCLIP can be high-quality proxies to enable the privacy-preserving analysis of proprietary data.
@@ Methods
We compare our method against competing methods, including CLIP, MedCLIP, PubMedCLIP, PLIP, Med-Flamingo and LLaVA-Med. MedCLIP [30] extends the pretraining to include large unpaired images and texts through contrastive learning. It uses the pretrained BioClinicalBERT and Swin Transformer [54] as the backbone text encoder and visual encoder respectively, and fine-tunes on MIMIC-CXR and CheXpert datasets. PubMedCLIP [29] fine-tunes CLIP on the Radiology Objects in COntext (ROCO) dataset [17], which consists of 80K radiology image-text pairs drawn from PubMed articles. PLIP [7] was pretrained on OpenPath, which contains about 208K pathology images paired with natural language descriptions from medical Twitter. Med-Flamingo [10] is a multimodal few-shot learner based on OpenFlamingo-9B, further pre-trained with 2M interleaved medical image-text data from publications and textbooks, formulated for open-ended visual question answering. LLaVA-Med [38] extends BiomedCLIP and is further tuned to master open-ended conversational semantics using GPT-4 generated instruction-following data. All the models, except the last two which use prompts to answer close-ended questions, are adapted into ELEVATER for evaluation.
@@
We utilize the METER [36] framework to facilitate our experiments on visual question answering (VQA). It formulates the VQA task as a classification task. The core module of METER is a transformer-based co-attention multimodal fusion module that produces cross-modal representations over the image and text encodings, which are then fed to a classifier for predicting the final answer. We compare BiomedCLIP with general-domain CLIP, MAML (Model-Agnostic Meta-Learning) network pretrained only on visual data and PubMedCLIP. All three models were fine-tuned for VQA tasks using the QCR (Question answering via Conditional Reasoning) framework [55] that alternatively uses a MLP-based attention networks with conditional reasoning as the fusion module. Results for MAML, CLIP and PubMedCLIP are collected from [29]. We also compare our method with instruction-aware fine-tuned models BiomedGPT-B [6], which was fine-tuned on 25 curated datasets, LLaVA-Med and significantly larger model Med-PaLM M (562B) [37], which was fine-tuned on 12 open source datasets and 14 individual tasks.
We evaluated our method and competing methods on VQA-RAD [56] and SLAKE [57] using accuracy and token-level F1 metrics. The accuracy metric evaluates string-level correctness by treating VQA as a classification task with a set of predefined answer candidates collected from the training data, while the token-level F1 metric assesses token-level matching by feeding the answers into GPT-2 to obtain tokens. The dataset VQA-RAD consists of $315$ radiology images and $3,515$ question-answer pairs that were manually constructed by clinicians. Images in the test set are also present in the training set but the question-answer pairs do not overlap. SLAKE (English only) consists of $642$ radiology images and over $7,000$ question-answer pairs annotated by experienced physicians. It covers more human body parts than VQA-RAD and does not have common images between the training and test sets. See Supplementary Table 1 for details.
Privacy-preserving proprietary data analysis
We collected 980 chest X-ray images and the associated clinical reports from Providence Healthcare organization and treated them as proprietary data. We employed BiomedCLIP for image-to-image retrieval to find the most similar images in the PMC-15M dataset. To optimize computational efficiency, we downsampled PMC-15M in two steps. Initially, we sampled $1,000$ MIMIC-CXR seed images and used BiomedCLIP to retrieve $100,000$ CXR-relevant images from PMC-15M. Subsequently, we utilized BiomedCLIP to filter out the 90% least consistent image-caption pairs, based on low similarity scores between image and caption embeddings, resulting in a refined subset of $10,000$ CXR images from PMC-15M used for final image-to-image retrieval.
We obtained 13 CheXbert labels, excluding ‘No Findings’, for the proprietary images by running CheXbert on the Findings and Impression sections of their corresponding clinical reports. Each label was classified as either positive or negative, with uncertain, blank, and negative outputs from CheXbert all categorized as negative. For the retrieved PMC images, we acquired labels by applying CheXbert to their corresponding captions. We conducted similar image-to-image retrieval experiments using PLIP and CLIP, evaluating label agreement between the top one retrieved and proprietary images through F1 and Recall metrics.
@@
where $\tau$ is a learnable temperature parameter, directly optimized during training as a log-parameterized multiplicative scalar; ${\bm{I}}_{i}$ and ${\bm{T}}_{i}$ are embeddings for the $i$-th image and text, produced by a linear projection layer on top of the image encoder and text encoder. Rather than initializing with pretrained weights, CLIP trains the image encoder and text encoder from scratch. For the image encoder, CLIP considers two different architectures, ResNet-50 [58] and Vision Transformer [ViT; 31]. The text encoder is effectively GPT-2 [28] based on transformer [59].
@@
On the text side, we replace a blank-slate GPT-2 with a pretrained language model more suited for biomedicine. Specifically, we initialize with PubMedBERT, which shows substantial gains from domain-specific pretraining [24]. Correspondingly, for the tokenizer, we replace Byte-Pair Encoding [BPE; 60] with WordPiece [61], which uses unigram-based likelihood rather than shattering all words to characters and greedily forming larger tokens based on frequency. The original CLIP uses a context of 77 tokens, but biomedical text is typically longer, as shown in Fig. 1A. We thus increase the context size to 256, which covers 90% of PMC captions. Supplementary Table 1 shows that both modifications bring substantial improvements over the original CLIP model on the validation set.